OpenStudy (anonymous):

get dy/ dx if this please. x^3 - 7x^2y^3 + 4y^2 = -16

OpenStudy (anonymous):

This is implicit differnetiation, so what we do is differntiatet both sides of the equation

OpenStudy (anonymous):

so, we get this: 3x^2-14x2y^3+6y^2y'7x^2=0

OpenStudy (anonymous):

Awww what the hey, I decided that if I sleep early I'll probably stay awake all night. If you want, I can show you how this is done. Wanna work on it together?

OpenStudy (anonymous):

ok. let's do it.

OpenStudy (anonymous):

Awesome. First thing we do is put each side in parentheses.

OpenStudy (anonymous):

(x^3 - 7x^2y^3 + 4y^2) = (-16)

OpenStudy (anonymous):

Then we write dy/dx next to each parentheses. dy/dx(x^3 - 7x^2y^3 + 4y^2) = dy/dx(-16) So far so good?

OpenStudy (anonymous):

yeah,

OpenStudy (anonymous):

Alright. Do you remember the rule for how you can differentiate each term separately if there's a plus or minus sign between them? Like how dy/dx (4x^2 + 3x) can be split into dy/dx(4x^2) + dy/dx(3x)

OpenStudy (anonymous):

(differentiate is what some people accidentally call "deriving", just in case you were wondering)

OpenStudy (anonymous):

Lemme know if that part makes sense so far, otherwise it'll be too confusing if we start the next steps.

OpenStudy (anonymous):

yeah, i'm still getting it.

OpenStudy (anonymous):

Awesome. So go ahead and tell me how we will rewrite the current equation, dy/dx(x^3 - 7x^2y^3 + 4y^2) = dy/dx(-16) when we separate it by parts into separate dy/dx parts.

OpenStudy (anonymous):

(Don't do any actual math on it yet. Just distribute the dy/dx)

OpenStudy (anonymous):

Again, if anything's confusing, please let me know before you bail out. For some reason lots of folks bail out instead of telling me that they're stuck. :P

OpenStudy (anonymous):

Did you give up? :(

OpenStudy (anonymous):

hey i'm still on it.

OpenStudy (anonymous):

Ah ok. I thought you left me. :P Remember: you just have to set it up, don't do any math yet.

OpenStudy (anonymous):

yes, i know

OpenStudy (anonymous):

Okie dokie. Just post it here when you've set it up.

OpenStudy (anonymous):

this is my answer. y' = (3x^2) / [2y (21x^2 - 4)]

OpenStudy (anonymous):

Okie dokie, lemme check it out. You did the math instead of just setting it up, so I wasn't quite ready to get an answer. Haha.

OpenStudy (anonymous):

sorry. i'm quite on a hurry.

OpenStudy (anonymous):

y' = (3x^2-14xy^3)/(7x^23y^2+8y) is what I got.

OpenStudy (anonymous):

i answered it wrong! i'm so sorry.

OpenStudy (anonymous):

let me try again please.

OpenStudy (anonymous):

Not a problem. Sure thing. But let's do it step by step. Just tell me what it looks like after this step: dy/dx(x^3 - 7x^2y^3 + 4y^2) = dy/dx(-16) Just show me where we put each dy/dx after we get rid of the ( )

OpenStudy (anonymous):

y' = (3x^2 - 14xy^3) / (21x^2y^2 - 8y) this is my answer. please kindly check it. :)

OpenStudy (anonymous):

That isn't what I got. :/ But let's just go through it one step at a time, that way we can correct each other if one of us makes a mistake. I'll go ahead and do the step that I've been begging you to try. dy/dx(x^3 - 7x^2y^3 + 4y^2) = dy/dx(-16) equals dy/dx(x^3) - dy/dx(7x^2y^3) + dy/dx(4y^2) = dy/dx(-16) Does this make sense so far, and does it look correct?

OpenStudy (anonymous):

geez. i really don't have any more time. my exam will be in about 20 minutes and i still have to go to school. i'm so sorry. but you're a good teacher so far. i hope you can teach me again next time. just we me luck. :)

OpenStudy (anonymous):

Sure thing. Good luck. Just one quick notice: dy/dx(7x^2y^3) On this step MAKE SURE YOU USE THE PRODUCT RULE! Just in case you haven't. And remember: dy/dx of y^2 is 2y*y'

OpenStudy (anonymous):

yep. i'll keep it in mind.. lots of thanks!

OpenStudy (anonymous):

and constants are automatic zero! :))

OpenStudy (anonymous):

Awesome! I think you'll do very well. Get back to me after the test if you have any questions!

OpenStudy (anonymous):

Kinda. Just one mistake: does 1 + 2 = 31?